# Gamma-ray Bursts as Distance Indicators by a Statistical Learning Approach

### v2

@@ 1 Introduction
As mentioned above, using correlations has allowed us to obtain pseudo-redshifts for larger samples for cosmological studies (Aldowma & Razzaque, 2024; Lloyd-Ronning, 2022; Atteia, 2003; Yonetoku et al., 2004; Kocevski & Liang, 2006; Tsutsui et al., 2011; Dainotti et al., 2011; Zhang & Wang, 2018), but the inferred redshifts are not yet accurate. We propose a very effective method using statistical learning tools for determining redshifts.
@@ 3.2 Luminosity function and density rate evolution
However, the values of these luminosity breaks are lower than those in Pescalli et al. (2016); Dong et al. (2022); Yu et al. (2015); Petrosian et al. (2015); Wanderman & Piran (2015, 2010), where the average log luminosity break is $51$. This is because we use plateau luminosity for deriving the LF. The difference between the $\log$ flux of the prompt and plateau is of the order of $5$. This estimate is taken as the difference between between an average of the logarithm of the peak flux of the prompt emission and the logarithm of flux at the end of the plateau emission in our sample. If we account for this in our luminosity break, then our values aligns with the literature.
@@ 5 Acknowledgement
We are particular grateful to Michal Ostrowski for his precious discussion during the data analysis and results, to Trevor Nelson for his preliminary analysis during the summer internship program at Stanford University supported by SLAC, to Simona Scimemi for the preliminary analysis related only to prompt emission variables during the winter internship 2018 at INAF Bologna, to Monica Brora on the very preliminary analysis on random forest. We also grateful to Macjei Bilicki for the useful discussion on the parameter space division. We thank particularly Douglas Figberg for his interesting discussion on selection biases related to this procedure and the way on how to solve them. M.G.D is particularly grateful for the support received as RIKEN pionieer research during her stay in Riken. M.G.D is grateful to the Gill Ferrandes and Donald Warren for their interesting discussions on the method of SuperLearner. We are also grateful to Trevor Hastie for useful suggestions to avoid overfitting and piece of the code for the 10-fold cross-validation. We are also grateful for the initial discussion on the project to Dmitry Maleshev. M.G.D. acknowledges the funding from MINIATURA2, Numer: 2018/02/X/ST9/03673. M.G.D also acknowledges Prof. Peter Behroozi for the insightful discussion on how to avoid overfitting in our statistical learning models. M.G.D and A.N also thank Nikita Khatiya, Dhruv Bal and Dr. Dieter Hartmann for extremely valuable feedback on the DRE analysis. This research was also support by the Visibility and Mobility module of the Jagiellonian University (Grant number: WSPR.WSDNSP.2.5.2022.5) and the NAWA STER Mobility Grant (Number: PPI/STE/2020/1/00029/U/00001). We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032. This work was also supported by the Polish National Science Centre grant UMO-2018/30/M/ST9/00757 and by Polish Ministry of Science and Higher Education grant DIR/WK/2018/12.
@@ A.1.1 Predictors selection
The GRB sample contains a total of nine predictors. Four belong to the prompt: $Fluence$, $T_{90}$, $\log(Peakflux)$, and $Photon$ $Index$. The rest belong to the plateau and afterglow: $T_{a}$, $F_{a}$, $\alpha$, $\beta$, and $NH$. To pick out the best predictors of redshift among these 9 variables, we applied the Least Absolute Shrinkage and Selection Operator (LASSO) predictor selection technique. The LASSO technique applies a shrinkage method for linear regression by forcing the $\ell^{1}$ norm (sum of the magnitude of all vectors in the given space) of the result vector to be $\leq\lambda$, where $\lambda$ is a positive number known as the tuning parameter. This form of penalization enables the model to choose specific predictors while nullifying others by assigning their coefficients a value of 0 (Tibshirani, 1996). The tuning parameter determines the degree of shrinkage applied to the estimated vector. Consequently, when using fewer predictors, the model becomes more interpretable and generally exhibits lower prediction errors than the full model. For our analysis, we use the GLMNET function with the LASSO selection predictors (Hastie et al., 2017; Tibshirani et al., 2012). We pick the $\lambda.1se$ value, the maximum $\lambda$ value for which the error is within one standard deviation (Friedman et al., 2010a), and its corresponding coefficients for the predictors s. Thus, we drop $T_{90}$ and $Fluence$ while keeping the rest of the 7 variables. The coefficients assigned by LASSO to each of them are displayed in Fig. 6, and we choose only the non-zero coefficient predictors. The LASSO selected predictors are also presented in Fig. 1.

### v4

@@ 1 Introduction
As mentioned above, using correlations has allowed us to obtain pseudo-redshifts for larger samples for cosmological studies (Lloyd-Ronning, 2022; Atteia, 2003; Yonetoku et al., 2004; Kocevski & Liang, 2006; Tsutsui et al., 2011; Dainotti et al., 2011; Zhang & Wang, 2018), but the inferred redshifts are not yet accurate. Recently, Aldowma & Razzaque (2024) trained a Deep Neural Networks to predict the redshift of Fermi-GBM and Konus Wind GRBs using spectral fit parameters. However, they do not perform cross-validation and do not eliminate duplicated GRBs from the two data sets. Regardless, although not reliable, their results do highlight the strength of such an approach. We propose a very effective method using statistical learning tools for determining redshifts.
@@ 3.2 Luminosity function and density rate evolution
However, the values of these luminosity breaks are lower than those in Pescalli et al. (2016); Dong et al. (2022); Yu et al. (2015); Petrosian et al. (2015); Wanderman & Piran (2015, 2010), where the average log luminosity break is $51$. This is because we use plateau luminosity for deriving the LF. The difference between the $\log$ flux of the prompt and plateau is of the order of $5$. This estimate is taken as the difference between an average of the logarithm of the peak flux of the prompt emission and the logarithm of flux at the end of the plateau emission in our sample. If we account for this in our luminosity break, then our values align with the literature.
@@ 5 Acknowledgement
We are particular grateful to Michal Ostrowski for his precious discussion during the data analysis and results, to Trevor Nelson for his preliminary analysis during the summer internship program at Stanford University supported by SLAC, to Simona Scimemi for the preliminary analysis related only to prompt emission variables during the winter internship 2018 at INAF Bologna, to Monica Brora on the very preliminary analysis on random forest. We also grateful to Macjei Bilicki for the useful discussion on the parameter space division. We thank particularly Douglas Figberg for his interesting discussion on selection biases related to this procedure and the way on how to solve them. M.G.D is particularly grateful for the support received as RIKEN pionieer research during her stay in Riken. M.G.D is grateful to the Gill Ferrandes and Donald Warren for their interesting discussions on the method of SuperLearner. We are also grateful to Trevor Hastie for useful suggestions to avoid overfitting and piece of the code for the 10-fold cross-validation. We are also grateful for the initial discussion on the project to Dmitry Maleshev. M.G.D. acknowledges the funding from MINIATURA2, Numer: 2018/02/X/ST9/03673. M.G.D also acknowledges Prof. Peter Behroozi for the insightful discussion on how to avoid overfitting in our statistical learning models. M.G.D and A.N also thank Nikita Khatiya, Dhruv Bal and Dr. Dieter Hartmann for extremely valuable feedback on the DRE analysis. This research was also support by the Visibility and Mobility module of the Jagiellonian University (Grant number: WSPR.WSDNSP.2.5.2022.5) and the NAWA STER Mobility Grant (Number: PPI/STE/2020/1/00029/U/00001). AN would also like to acknowledge the NAOJ Exploratory Grant and Division of Science for supporting his stay in Tokyo. We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032. This work was also supported by the Polish National Science Centre grant UMO-2018/30/M/ST9/00757 and by Polish Ministry of Science and Higher Education grant DIR/WK/2018/12. Numerical computations were in part carried out on Small Parallel Computers at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.
@@ A.1.1 Predictors selection
The GRB sample contains a total of nine predictors. Four belong to the prompt: $Fluence$, $T_{90}$, $\log(Peakflux)$, and $Photon$ $Index$. The rest belong to the plateau and afterglow: $T_{a}$, $F_{a}$, $\alpha$, $\beta$, and $NH$. To pick out the best predictors of redshift among these 9 variables, we applied the Least Absolute Shrinkage and Selection Operator (LASSO) predictor selection technique. The LASSO technique applies a shrinkage method for linear regression by forcing the $\ell^{1}$ norm (sum of the magnitude of all vectors in the given space) of the result vector to be $\leq\lambda$, where $\lambda$ is a positive number known as the tuning parameter. This form of penalization enables the model to choose specific predictors while nullifying others by assigning their coefficients a value of 0 (Tibshirani, 1996). The tuning parameter determines the degree of shrinkage applied to the estimated vector. Consequently, when using fewer predictors, the model becomes more interpretable and generally exhibits lower prediction errors than the full model. For our analysis, we use the GLMNET function with the LASSO selection predictors (Hastie et al., 2017; Tibshirani et al., 2012). We pick the $\lambda.1se$ value, the maximum $\lambda$ value for which the error is within one standard deviation (Friedman et al., 2010a), and its corresponding coefficients for the predictors. Thus, we drop $T_{90}$ and $Fluence$ while keeping the rest of the 7 variables. The coefficients assigned by LASSO to each of them are displayed in Fig. 6, and we choose only the non-zero coefficient predictors. The LASSO selected predictors are also presented in Fig. 1.